# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 6.1 Cohomological descent from $G/B$ to $B$
Let $\mathcal{F}=\mathcal{L}({\sf F})={G}\times^{B}\sf F$ be a $G$-equivariant sheaf on $G/B$ given by a finitely generated $B$-module $\sf F$, which is projective over ${\sf k}$. By [29, I 3.3, Proposition 5.12] there are isomorphisms
${\rm H}^{p}({B},{\sf k}[{G}]\otimes{\sf F})=\mathrm{R}^{p}{\mathrm{ind}}_{B}^{G}(\sf F)={\rm H}^{p}({G/B},\mathcal{F}),$ (4)
@@
Assume $H^{p}(B,{\sf F}\otimes\nabla_{\mu})=0$ for $p>0$ and all $\mu$.
@@
    Then ${\rm H}^{p}({G/B},\mathcal{F})=0$ for all $p>0$ and ${\rm H}^{0}({G/B},\mathcal{F})$ has a good filtration with with associated graded
    $\bigoplus_{\lambda\in X({T})_{+}}\nabla_{\lambda}\otimes{\rm H}^{0}({B},\nabla_{-w_{0}\lambda}\otimes{\sf F}).$
@@
    If moreover $H^{0}(B,{\sf F}\otimes\nabla_{\mu})=0$ for all $\mu$, then ${\rm H}^{0}({G/B},\mathcal{F})=0$.
@@
    $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(\lambda)\otimes Q(\mu))=0$ for $i>0$.
@@
    ${\mathrm{ind}}_{B}^{G}(P(\lambda)\otimes Q(\mu))$ has a good filtration.
@@
$\langle[\mathcal{L}(P(\lambda))],[\mathcal{L}(Q(\mu))]\rangle={\mathrm{ind}}_{B}^{G}(P(\lambda)\otimes Q(\mu))$.∎
The proof of Theorem 6.2 also works over ${\mathbb{Z}}$. Or one may use the universal coefficient theorem [29, I 4.18]. Thus
@@
    ${\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))$ vanishes unless $w\leq v$.
@@
    ${\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{v}))={\sf k}$.
@@
The ${\mathrm{ind}}_{\mathbb{B}}^{\mathbb{G}}(P(\lambda)_{\mathbb{Z}}\otimes Q(\mu)_{\mathbb{Z}})$ are free ${\mathbb{Z}}$-modules of finite rank, so the universal coefficient theorem [29, I 4.18] shows that the theorem generalizes from ${\sf k}={\mathbb{C}}$ to ${\mathbb{Z}}$, which in turn implies the case where ${\sf k}$ is a field, again by the universal coefficient theorem.
@@
    $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{v}))=\begin{cases}{\sf k}&\text{if $i=0$}\\ 0&\text{else.}\end{cases}$
@@
    If $w\not\leq v$ then $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))=0$ for all $i$.
@@
Let $M$, $N$ be finite dimensional $G$-modules.
@@
    ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{v}))={\mathrm{Ext}}^{i}_{G}(M,N)$ for all $i$.
@@
    If $w\not\leq v$ then ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{w}))=0$ for all $i$.
@@
$(1).$ We have a spectral sequence [29, I Proposition 4.5]
${\mathrm{Ext}}^{i}_{G}(M\otimes N^{*},\mathrm{R}^{j}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w})))\Rightarrow{\mathrm{Ext}}^{i+j}_{B}(M\otimes N^{*},P(-e_{v})\otimes Q(e_{w})).$
By Theorem 6.2 (1), $\mathrm{R}^{j}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))=0$ for $j>0$. Thus, the above spectral sequence degenerates, and with Theorem 6.6 this gives ${\mathrm{Ext}}^{i}_{B}(M\otimes N^{*},P(-e_{v})\otimes Q(e_{v}))={\mathrm{Ext}}^{i}_{G}(M\otimes N^{*},\mathrm{R}^{0}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{v})))={\mathrm{Ext}}^{i}_{G}(M\otimes N^{*},{\sf k})={\mathrm{Ext}}^{i}_{G}(M,N)$.
$(2).$ If $w\not\leq v$, the above spectral sequence still degenerates by Theorem 6.2, and Theorem 6.6 gives ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{w}))=0$ for all $i$. ∎
@@ 7.1 Triangulated hull
Definition 10 (Triangulated hull).
@@ 7.2 Thick hull
Definition 12 (Thick hull).
@@ 7.3 Admissible subcategories
We follow the exposition of [20, Section 2.1]. Let $\mathcal{N}\subset\sf D$ be a full triangulated subcategory. The right orthogonal to $\mathcal{N}$ is the full subcategory $\mathcal{N}^{\perp}\subset\sf D$ consisting of all objects $X$ such that $\mathop{\mathrm{Hom}}\nolimits_{\sf D}(Y,X)=0$ for any $Y\subset\mathcal{N}$. The left orthogonal ${}^{\perp}\mathcal{N}$ is defined analogously. The orthogonals are also triangulated subcategories.
Definition 13.
@@
[6, Lemma 3.1] Let $\mathcal{N}\subset\mathcal{D}$ be a full triangulated subcategory, and consider its right orthogonal $\mathcal{N}^{\perp}\subset\mathcal{D}$. Then the following are equivalent:
@@
[30, Proposition 3.2.8] Let $\mathcal{N}\subset\mathcal{D}$ be a thick subcategory. Then the following are equivalent:
@@ 7.4 Semiorthogonal decompositions
Definition 14.
A semiorthogonal decomposition of a triangulated category $\mathcal{D}$ is a sequence of full triangulated subcategories $(\mathcal{N}_{1},\dots,\mathcal{N}_{n})$ in $\mathcal{D}$ such that $\mathcal{N}_{i}\subseteq\mathcal{N}_{j}^{\perp}$ for $1\leq i<j\leq n$ and for every object $X\in\mathcal{D}$ there exists a chain of morphisms in $\mathcal{D}$,
@@ 7.5 Mutations
[32, Lemma 2.7] There are equivalences ${\bf L}_{{\sf A}}:{\sf B}\simeq\mathcal{D}/{\sf A}\simeq{\bf L}_{{\sf A}}({\sf B})$ and ${\bf R}_{{\sf A}}:{\sf A}\simeq\mathcal{D}/{\sf B}\simeq{\bf R}_{{\sf A}}({\sf B})$.
Definition 16.
@@
[32, Lemma 2.10] Let $\mathcal{D}=\langle{\sf A}_{1},\dots,{\sf A}_{n}\rangle$ be a semiorthogonal decomposition such that the components ${\sf A}_{k}$ and ${\sf A}_{k+1}$ are completely orthogonal, i.e., $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\sf A}_{k},{\sf A}_{k+1})=0$ and $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\sf A}_{k+1},{\sf A}_{k})=0$. Then
@@ 7.7 Admissible subcategories from exceptional collections
One proves [6, Theorem 3.2] that the subcategory ${\mathrm{hull}}(E_{0},\dots,E_{n})\subseteq\mathcal{D}$ is admissible in $\mathcal{D}$. If the collection is full, then ${\mathcal{D}}={\mathrm{hull}}(E_{0},\dots,E_{n})$.
@@
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(E,F)\otimes E\rightarrow F\rightarrow{\bf L}_{\langle E\rangle}(F),\qquad{\bf R}_{\langle E\rangle}(F)\rightarrow F\rightarrow\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(F,E)^{\ast}\otimes E.$ (13)
@@ 7.8 Dual exceptional collections
[20, Proposition 2.15] Let $(E_{0},\dots,E_{n})$ be a full exceptional collection in a triangulated category $\mathcal{D}$. The left dual exceptional collection $\langle F_{n},\dots,F_{0}\rangle$ is uniquely determined by the following property:
@@
8 Generating the categories ${\mathrm{rep}}(B)$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$
8.1 The categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$
@@
We recall that ${\mathrm{Rep}}(B)$ (resp., ${\mathrm{rep}}(B)$) denotes the abelian category of rational $B$-modules (resp., the abelian subcategory of rational $B$-modules which are finitely generated over ${\sf k}$). We let denote $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(B))$ (resp., $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(G))$) the unbounded derived category of ${\mathrm{Rep}}(B)$ (resp., of ${\mathrm{Rep}}(G)$), and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ (resp., $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$) the bounded derived category of the small category ${\mathrm{rep}}(B)$ (resp., of ${\mathrm{rep}}(G)$).
Let ${\mathrm{rep}}_{\mathrm{fr}}(B)$ denote the full subcategory of ${\mathrm{rep}}(B)$ consisting of the representations that are free over ${\sf k}$. Define ${\mathrm{rep}}_{\mathrm{fr}}(G)$ similarly.
@@
$-\;\otimes_{\sf k}-:{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B)$
and (the internal $\mathop{\mathrm{Hom}}\nolimits_{\sf k}$ on ${\mathrm{rep}}_{\mathrm{fr}}(B)$):
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B).$
Put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$, for $M\in{\mathrm{rep}}_{\mathrm{fr}}(B)$. By [29, I, Section 2.7, (4)], there is a natural $B$-module structure on $M^{*}$.
@@
This is a special case of [43, Proposition 3]. ∎
@@
The Proposition implies in standard fashion that every bounded complex in ${\mathrm{rep}}(B)$ is quasi-isomorphic to a bounded complex with objects in ${\mathrm{rep}}_{\mathrm{fr}}(B)$. The above functors can thus be extended as exact bifunctors to the bounded derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$:
$(-)\otimes_{\sf k}(-):\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\otimes\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$
@@
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\times\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\to\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B)).$
In other words, we are using that by the dual of [9, Theorem 10.22, Remark 10.23], the bounded derived category of the abelian category ${\mathrm{rep}}(B)$ is equivalent to the bounded derived category of its exact subcategory ${\mathrm{rep}}_{\mathrm{fr}}(B)$.
We often write $\otimes_{\sf k}$ as $\otimes$. We put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$. By [29, I, Section 2.7, (5)], there is an isomorphism $\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,N)=M^{\ast}\otimes N$ for $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. There is $(\mathop{\mathrm{Hom}}\nolimits_{\sf k},\otimes_{\sf k})$-adjunction:
@@
Let $\mathop{\mathrm{D}\kern 0.0pt}\nolimits_{{\mathrm{rep}}(B)}({\mathrm{Rep}}(B))$ denote the derived category of complexes of ${\mathrm{Rep}}(B)$ whose cohomology lie in ${\mathrm{rep}}(B)$.
@@
There is an equivalence of triangulated categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{{\mathrm{rep}}(B)}({\mathrm{Rep}}(B))$.
@@
This is a particular case of [37, Appendix, Lemma A.3]. ∎

### v3

@@ 6.1 Cohomological descent from $\mathbf{G}/\mathbf{B}$ to $\mathbf{B}$
Let $\mathcal{F}=\mathcal{L}({\sf F})={\mathbf{G}}\times^{\mathbf{B}}\sf F$ be a $\mathbf{G}$-equivariant sheaf on $\mathbf{G}/\mathbf{B}$ given by a finitely generated $\mathbf{B}$-module $\sf F$, which is projective over ${\sf k}$. By [32, I 3.3, Proposition 5.12] there are isomorphisms
${\rm H}^{p}({\mathbf{B}},{\sf k}[{\mathbf{G}}]\otimes{\sf F})=\mathrm{R}^{p}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(\sf F)={\rm H}^{p}({\mathbf{G}/\mathbf{B}},\mathcal{F}),$ (7)
@@
Assume $H^{p}(\mathbf{B},{\sf F}\otimes\nabla_{\mu})=0$ for $p>0$ and all $\mu$.
@@
    Then ${\rm H}^{p}({\mathbf{G}/\mathbf{B}},\mathcal{F})=0$ for all $p>0$ and ${\rm H}^{0}({\mathbf{G}/\mathbf{B}},\mathcal{F})$ has a good filtration with with associated graded
    $\bigoplus_{\lambda\in X({\mathbf{T}})_{+}}\nabla_{\lambda}\otimes{\rm H}^{0}({\mathbf{B}},\nabla_{-w_{0}\lambda}\otimes{\sf F}).$
@@
    If moreover $H^{0}(\mathbf{B},{\sf F}\otimes\nabla_{\mu})=0$ for all $\mu$, then ${\rm H}^{0}({\mathbf{G}/\mathbf{B}},\mathcal{F})=0$.
@@
    $\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(\lambda)\otimes Q(\mu))=0$ for $i>0$.
@@
    ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(\lambda)\otimes Q(\mu))$ has a good filtration.
@@
$\langle[\mathcal{L}(P(\lambda))],[\mathcal{L}(Q(\mu))]\rangle={\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(\lambda)\otimes Q(\mu))$.∎
The proof of Theorem 6.2 also works over ${\mathbb{Z}}$. Or one may use the Universal coefficient Theorem 1.8. Thus
@@
    ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{w}))$ vanishes unless $w\leq v$.
@@
    ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{v}))={\sf k}$.
@@
The ${\mathrm{ind}}_{\mathbb{B}}^{\mathbb{G}}(P(\lambda)_{\mathbb{Z}}\otimes Q(\mu)_{\mathbb{Z}})$ are free ${\mathbb{Z}}$-modules of finite rank, so the Universal coefficient Theorem 1.8 shows that the theorem generalizes from ${\sf k}={\mathbb{C}}$ to ${\mathbb{Z}}$, which in turn implies the case where ${\sf k}$ is a field, again by the Universal coefficient Theorem.
@@
    $\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{v}))=\begin{cases}{\sf k}&\text{if $i=0$}\\ 0&\text{else.}\end{cases}$
@@
    If $w\not\leq v$ then $\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{w}))=0$ for all $i$.
@@
Let $M$, $N$ be finite dimensional $\mathbf{G}$-modules.
@@
    ${\mathrm{Ext}}^{i}_{\mathbf{B}}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{v}))={\mathrm{Ext}}^{i}_{\mathbf{G}}(M,N)$ for all $i$.
@@
    If $w\succ v$ then ${\mathrm{Ext}}^{i}_{\mathbf{B}}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{w}))=0$ for all $i$.
@@
$(1).$ We have a spectral sequence [32, I Proposition 4.5]
${\mathrm{Ext}}^{i}_{\mathbf{G}}(M\otimes N^{*},\mathrm{R}^{j}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{w})))\Rightarrow{\mathrm{Ext}}^{i+j}_{\mathbf{B}}(M\otimes N^{*},P(-e_{v})\otimes Q(e_{w})).$
By Theorem 6.2 (1), $\mathrm{R}^{j}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{w}))=0$ for $j>0$. Thus, the above spectral sequence degenerates, and with Theorem 6.6 this gives ${\mathrm{Ext}}^{i}_{\mathbf{B}}(M\otimes N^{*},P(-e_{v})\otimes Q(e_{v}))={\mathrm{Ext}}^{i}_{\mathbf{G}}(M\otimes N^{*},\mathrm{R}^{0}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{v})))={\mathrm{Ext}}^{i}_{\mathbf{G}}(M\otimes N^{*},{\sf k})={\mathrm{Ext}}^{i}_{\mathbf{G}}(M,N)$.
$(2).$ If $w\succ v$ then $w\not\leq v$, and the above spectral sequence still degenerates by Theorem 6.2, and Theorem 6.6 gives ${\mathrm{Ext}}^{i}_{\mathbf{B}}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{w}))=0$ for all $i$. ∎
@@ 7.1 Triangulated hull
Definition 7.1 (Triangulated hull).
@@ 7.2 Thick hull
Definition 7.3 (Thick hull).
@@ 7.3 Admissible subcategories
We follow the exposition of [23, Section 2.1]. Let $\mathcal{N}\subset\sf D$ be a full triangulated subcategory. The right orthogonal to $\mathcal{N}$ is the full subcategory $\mathcal{N}^{\perp}\subset\sf D$ consisting of all objects $X$ such that $\mathop{\mathrm{Hom}}\nolimits_{\sf D}(Y,X)=0$ for any $Y\in\mathcal{N}$. The left orthogonal ${}^{\perp}\mathcal{N}$ is defined analogously. The orthogonals are also triangulated subcategories.
Definition 7.4.
@@
[9, Lemma 3.1] Let $\mathcal{N}\subset\mathcal{D}$ be a full triangulated subcategory, and consider its right orthogonal $\mathcal{N}^{\perp}\subset\mathcal{D}$. Then the following are equivalent:
@@
[33, Proposition 3.2.8] Let $\mathcal{N}\subset\mathcal{D}$ be a thick subcategory. Then the following are equivalent:
@@ 7.4 Semiorthogonal decompositions
Definition 7.7.
A semiorthogonal decomposition of a triangulated category $\mathcal{D}$ is a sequence of full triangulated subcategories $(\mathcal{N}_{1},\dots,\mathcal{N}_{n})$ in $\mathcal{D}$ such that $\mathcal{N}_{i}\subset\mathcal{N}_{j}^{\perp}$ for $1\leq i<j\leq n$ and for every object $X\in\mathcal{D}$ there exists a chain of morphisms in $\mathcal{D}$,
@@ 7.5 Mutations
[35, Lemma 2.7] There are equivalences ${\bf L}_{{\sf A}}:{\sf B}\simeq\mathcal{D}/{\sf A}\simeq{\bf L}_{{\sf A}}({\sf B})$ and ${\bf R}_{{\sf A}}:{\sf A}\simeq\mathcal{D}/{\sf B}\simeq{\bf R}_{{\sf A}}({\sf B})$.
Definition 7.10.
@@
[35, Lemma 2.10] Let $\mathcal{D}=\langle{\sf A}_{1},\dots,{\sf A}_{n}\rangle$ be a semiorthogonal decomposition such that the components ${\sf A}_{k}$ and ${\sf A}_{k+1}$ are completely orthogonal, i.e., $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\sf A}_{k},{\sf A}_{k+1})=0$ and $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\sf A}_{k+1},{\sf A}_{k})=0$. Then
@@ 7.7 Admissible subcategories from exceptional collections
One proves [9, Theorem 3.2] that the subcategory ${\mathrm{hull}}(E_{0},\dots,E_{n})\subset\mathcal{D}$ is admissible in $\mathcal{D}$. If the collection is full, then ${\mathcal{D}}={\mathrm{hull}}(E_{0},\dots,E_{n})$.
@@
$\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(E,F)\otimes E\rightarrow F\rightarrow{\bf L}_{\langle E\rangle}(F),\qquad{\bf R}_{\langle E\rangle}(F)\rightarrow F\rightarrow\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(F,E)^{\ast}\otimes E.$ (16)
@@ 7.8 Dual exceptional collections
[23, Proposition 2.15] Let $(E_{0},\dots,E_{n})$ be a full exceptional collection in a triangulated category $\mathcal{D}$. The left dual exceptional collection $\langle F_{n},\dots,F_{0}\rangle$ is uniquely determined by the following property:
@@
8 Generating the categories ${\mathrm{rep}}({\mathbf{B}})$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$
8.1 The categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$
@@
We recall that ${\mathrm{Rep}}(\mathbf{B})$ (resp., ${\mathrm{rep}}(\mathbf{B})$) denotes the abelian category of rational $\mathbf{B}$-modules (resp., the abelian subcategory of rational $\mathbf{B}$-modules which are finitely generated over ${\sf k}$). We let denote $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}({\mathbf{B}}))$ (resp., $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(\mathbf{G}))$) the unbounded derived category of ${\mathrm{Rep}}(\mathbf{B})$ (resp., of ${\mathrm{Rep}}(\mathbf{G})$), and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$ (resp., $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$) the bounded derived category of the smaller category ${\mathrm{rep}}(\mathbf{B})$ (resp., of ${\mathrm{rep}}(\mathbf{G})$).
Let ${\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$ denote the full subcategory of ${\mathrm{rep}}(\mathbf{B})$ consisting of the representations that are free over ${\sf k}$. Define ${\mathrm{rep}}_{\mathrm{fr}}(\mathbf{G})$ similarly.
@@
$-\;\otimes_{\sf k}-:{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\times{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\to{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$
and (the internal $\mathop{\mathrm{Hom}}\nolimits_{\sf k}$ on ${\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$):
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\times{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\to{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B}).$
Put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$, for $M\in{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$. By [32, I, Section 2.7, (4)], there is a natural $\mathbf{B}$-module structure on $M^{*}$.
@@
This is a special case of [45, Proposition 3]. ∎
@@
The Proposition implies in standard fashion that every bounded complex in ${\mathrm{rep}}(\mathbf{B})$ is quasi-isomorphic to a bounded complex with objects in ${\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$. The above functors can thus be extended as exact bifunctors to the bounded derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$:
$(-)\otimes_{\sf k}(-):\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))\otimes\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$
@@
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))\times\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))\to\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B})).$
In other words, we are using that by the dual of [12, Theorem 10.22, Remark 10.23], the bounded derived category of the abelian category ${\mathrm{rep}}(\mathbf{B})$ is equivalent to the bounded derived category of its exact subcategory ${\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$.
We often write $\otimes_{\sf k}$ as $\otimes$. We put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$. By [32, I, Section 2.7, (5)], there is an isomorphism $\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,N)=M^{\ast}\otimes N$ for $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. There is $(\mathop{\mathrm{Hom}}\nolimits_{\sf k},\otimes_{\sf k})$-adjunction:
@@
Let $\mathop{\mathrm{D}\kern 0.0pt}\nolimits_{{\mathrm{rep}}({\mathbf{B}})}({\mathrm{Rep}}(\mathbf{B}))$ denote the derived category of complexes of ${\mathrm{Rep}}(\mathbf{B})$ whose cohomology lie in ${\mathrm{rep}}({\mathbf{B}})$.
@@
There is an equivalence of triangulated categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{{\mathrm{rep}}({\mathbf{B}})}({\mathrm{Rep}}(\mathbf{B}))$.
@@
This is a particular case of [40, Appendix, Lemma A.3]. ∎